We are to use interval notation for the solution set, I am still confused with this. Can you help?
2<2x+4<10
Here is another
3x-6=-3(2-x)
thankyou, lackinmathe

Answer by 303795(562)

2<2x+4<10
Subtract 4 from each part of the statement
2-4<2x+4-4<10-4
-2 < 2x < 6
ie -1 < x < 3

3x-6=-3(2-x)
3x-6=-6+3x
3x-6=3x-6
3x=3x
x=x
which is true for any value of x you would like to chose.